OpenStudy (anonymous):

solve compound inequality and graph 15≤(20+11+k)/3≤19

OpenStudy (gorv):

multiply each by 3

OpenStudy (gorv):

15*3<=20+11+k<=19*3 45<=31+K<=57

OpenStudy (gorv):

no subtract 31 from each